I need help on how to figure out problems like these:
-4=x-4
x-7=5( o already know this one but i need the steps and im not sure how to do it)
-2x+1=-5

Answer by funmath(2873)

Get the x all by itself on one side of the equal sign within the rules of algebra, whatever you do to one side of the equation, you have to do to the other side of the equation.
-4=x-4
-4+4=x-4+4
0=x+0
x=0
:
Check by letting x=0 and see if both sides =.
-4=(0)-4
-4=-4 We're right!!!!
:
x-7=5
x-7+7=5+7
x=12
:
Check
(12)-7=5
5=5 we're right again!!!
:
-2x+1=-5
-2x+1-1=-5-1
-2x=-6
-2x/-2=-6/-2
x=3
:
Check
-2(3)+1=-5
-6+1=-5
-5=-5 Right again!!!
Happy Calculating!!!!